The five term of a series A.P is 19 and 10th term is 39. Find the sum of 12th terms.
ย 

Grade:11

1 Answers

Arun
4 years ago
Dear Bipin
 
5th term T5 = 19 = a + 4 d
While
10th term = 39 = a + 9d
Now
Solve both the equations
Hence d = 4
So a = 3
Hence sum of 12 terms = (12/2)(2a + 11 d)
 = 6 (6 + 44) = 300
 
Regards
Arun